OpenStudy (anonymous):

log(sub4)x + log(sub4)(x-3) = 1?

OpenStudy (anonymous):

\[\log _{4}x + \log _{4}(x-3) = 1 \]

OpenStudy (anonymous):

If you "add" two logs you actually multiply them and if you "subtract" two logs you actually divide them. so x(x-3)=1. x^2-3x-1=0. solve the quadratic equation.